# RIVCoin: an alternative, integrated, CeFi/DeFi-Vaulted Cryptocurrency

### v1

@@ 6.1 Introduction
    Increasing the number of both informed and uninformed agents will reduce the negative impact played by noise traders.
@@
    After an increase in the number of agents, the return will be lower, as the result of the improved ability of prices in conveying information to the market, and the volatility will be lower as well.
  4. 4.
@@
As a general comment, increasing the number of agents improve the quality of the market as a whole, lowering price volatility. Increasing the number of agents creates a problem relative to the supply of the risky security: to cope with higher demand, without incurring into the risk to have price jumps, the direct consequence of this will be a cheaper price for the risky security. To guarantee a larger access to risky investments, making the demand higher, the single price should be smaller than in the case with a restricted demand.
Larger is the number of participants to the market, smaller is going to be the unitary price of each security, lower its volatility and higher the informational efficiency conveyed by prices.
@@ 6.2 The Model
Let us start with designing a market with asymmetric information. Assume there exists two groups of agents, both are assumed to be risk-averse: $N$ are assumed to be informed and $M$ are assumed to be uninformed. Intuitively, we can imagine that informed agents are professional traders and uninformed traders are individual non-professional traders, considering trading activity almost like a spare-time job. Furthermore, we assume the presence of noise traders in measure equal to $Z$: these individuals are assumed to be interpreted like automated traders.
@@ 6.2.1 Naïve Expectations
According to the naïve equilibrium outlined in Theorem 1, each trader does not act strategically, in the sense that when submitting an order, it is not taken into account the impact of her order on the equilibrium price or the market equilibrium.
@@ 6.3 Rational Expectations Equilibrium
On the other hand, informed agent use also the signal together with the current price to make expectations about future price, so that their expectation is $E(\tilde{S}|\tilde{X},p)$. If price is correlated with insiders’ signal and with the true value of the security, we will have that: $E(\tilde{S}|\tilde{X},p)=E(\tilde{S}|\tilde{X})$, and $Var(\tilde{S}|\tilde{X},p)=Var(\tilde{S}|\tilde{X})$.
@@ 6.5 Concluding Remarks
In this note by taking advantage of the results contained in Kyle (1989), DeJong and Rindi (2009) and Rindi (2008) we study the role of equilibrium asset pricing when the market is populated by group of agents with differential information, with different sizes. We considered a naïve equilibrium without any update of the beliefs and a fully revealing rational expectations equilibrium when non-informed agents can update their beliefs based on current prices.
In all cases the number of agents - both informed and uninformed - is crucial to determine the characteristics of the equilibrium. As the number of agents increases, prices increase their informative efficiency, enhancing the spreading out of the private information to the market. The Volatility of security’s price lowers. The return on the security lowers since agents need less protection against adverse price fluctuations.
These results show that by making the market for a security larger, open to a larger group of investors, improves the information conveyed by prices and makes private information to become publicly available. At the same time, the volatility lowers and the overall quality of the market information improves. The return will be lower, but this is the price that each agent will be willing to pay to have a market more open, transparent, liquid and less volatile.
A larger number of agents, given a fixed supply of the security, implies that each share will be issued to a lower price than in a context with a lower demand (because of a smaller number of traders).
As a consequence, a large number of players implies also a cheaper price, with lower volatility and better quality of information. The advantage of this situation is given by the fact that an higher number of individuals can be part of the market.
7 Roadmap
We plan to launch the cryptocurrency minting from British Virgin Islands, but already compliant with the main regulatory licenses in Middle East and Europe.
The Roadmap is divided into three main steps. It aims to build a fully functional product that will serve the entire RIV ecosystem as a foundational block, allowing access to an entire ecosystem of value.
Below you can find a detailed description of all the phases that will compose the launch of the RIVCoin. The timeline will be further expanded in the subsequent years when we plan to build even more added-value services for our Users.
@@ 7 Roadmap
    Phase 1: the Presale Phase where everyone interested in the project will be able to reserve a spot, purchasing batches of coins. Downloading RIV Wallet and completing the related KYC will allow the User to see the coins in the digital wallet, without any possibility to trade them yet.
@@
    Phase 2: the RIV/OSMO and/or USDC Liquidity Pool will be up and running. Users will be able to buy and sell RIVCoin on the secondary market. Also, Users will be able, in a subsequent phase, to bridge their tokens on Ethereum using Cøsmos bridges.
@@
    Phase 3: the Primary and the Secondary Market are both active. RIV DAO opens or closes Primary Minting according to market conditions and/or to the reserve’s appreciation, to stimulate or deter demand and with the sole objective of stabilizing the price around the reference value, until the max cap has been reached.
RIV Wallet will be further expanded by incrementally implementing the RIV Ecosystem. Detailed information will be provided in the following versions of this Whitepaper.
8 RIV Capital Group
Founded in 2021 and headquartered in Luxembourg [20], we focus on long-term investment strategies developed by carefully studying and analyzing investment opportunities.
At a proprietary and RAIF level, managed by an asset management company authorized and regulated, we deploy quantitative, model-driven strategies to develop insights about asset value based on statistical and fundamental analysis. Fuzzy Logic is used to combine and weight both the fundamental and statistical indicators with limited leverage.
Strategies involve investing in high-yield Fixed Income securities, focused on private debt (i.e., mini-bonds) and corporate bonds, combined with thematic equity portfolios based on the mega-trends characterizing the current economic wave. Our people are top investment bankers and traders with years of experience in the business and an undisputed reputation. A team of talented quant developers works and helps to implement and improve 24/7 the strategies and the automated system. We have profound knowledge of financial markets, blockchain, and DeFi. Passion, commitment, and competence are the guiding principles that sustain our growth.
The solution proposed is specifically thought for the Users. They have been traditionally subject to scams, frauds, financial illusions and fake hopes to become rich quickly.
This cryptocurrency is explicitly designed to offer a secured model of crypto and at the same time to be indeed a currency (medium of payment, fractionable, with a guarantor seller of last resort, etc.) starting from the first day. On top of that, when there is a Central Bank printing money we have to ask ourselves questions about inflation and seigniorage. RIVCoin is a cryptocurrency, but the decentralized seller of last resort will not create inflation and will democratically redistribute the seigniorage. This is very important to feed the reputation and the credibility of the cryptocurrency, as the scarcity is defined and the production of the reserves are clearly conveyed following an automated rule through the mechanisms of the DAO.
We could also open another chapter which is the skilled and professional use of the reserves which are active to produce returns through methods available to the top notch of the markets. It is with patience and a realistic aim that one can save money, accumulate returns and see personal wealth increasing in time.
The Founders invest themselves in this cryptocurrency and they expect both the strongest safety and also the best compounding in the time derived by the sane management of the reserves behind.
The User of any currency would like to base the use of it on a few factors: trust of the issuer, undisputed intrinsic value of the coin (i.e., Gold Standard until 1973) and accepted medium of exchange. The first and the third are a matter of gradual achievement. The second one is a matter of numbers. If one of the three comes less, then inefficiencies arise.
This is indeed a new paradigm: the integration of CeFi and DeFi, the stabilized cryptocurrency, the vaulted reserves and the related generated returns distributed to Liquidity Providers.

### v2

@@ 3.1 Information, Noise and Equilibrium
    Increasing the number of both informed and uninformed agents will reduce the negative impact played by noise traders: the return will be lower, as a result of the improved ability of prices to convey information to the market, but the volatility will be lower as well.
@@
As a general comment, increasing the number of agents improves the quality of the market as a whole, lowering price volatility and increasing the price level, converging to the equilibrium level. In a situation characterized by higher informational efficiency, security prices tend to increase. In this context, the role of both the fundamentals of the asset and its supply becomes crucial in order to balance the attractiveness of the security and to guarantee the access to a large number of individual investors. The important point to be stressed is that more agent buy the asset, higher is its price and lower its volatility for a fixed asset supply.
@@ 3.2 The Model
Let us start with designing a market with asymmetric information. Assume there exists two groups of agents, both are assumed to be risk-averse: $N$ are assumed to be informed and $M$ are assumed to be uninformed. Intuitively, we can imagine that informed agents are professional traders and uninformed agents are non-professional traders, considering trading activity almost like a spare-time job. Furthermore, we assume the presence of noise traders or agents in measure equal to $Z$: these individuals are assumed to be interpreted like automated traders.
@@ 3.2.1 Naïve Expectations
According to the naïve equilibrium outlined in Theorem 1, each trader does not act strategically, in the sense that when submitting an order, she does not take into account the impact of her order on the equilibrium price or market equilibrium.
@@ 3.3 Rational Expectations Equilibrium
On the other hand, informed agent adopt the signal together with the current price in order to make expectations about future price, so that their expectation is $E(\tilde{S}|\tilde{X},p)$. If price is correlated with insiders’ signal and with the true value of the security, we have that: $E(\tilde{S}|\tilde{X},p)=E(\tilde{S}|\tilde{X})$, and $Var(\tilde{S}|\tilde{X},p)=Var(\tilde{S}|\tilde{X})$.
@@ 3.5 Concluding Remarks
Building on insights from [kyle89, rindi08, djrindi], we have examined how varying the number of informed and uninformed participants influences equilibrium asset pricing. In both naïve scenarios without belief updates and more sophisticated frameworks where uninformed agents refine their assessments based on current prices, a key finding remains the same: increasing the number of participants amplifies the market’s informative efficiency and lowers price volatility. Although larger markets tend to have lower returns–reflecting a reduced need to compensate traders for potential adverse price movements–they benefit from broader liquidity, more robust price discovery, and greater collective access to information.
By attracting a wide spectrum of users, RIV Coin leverages these principles to establish a stable reserve in the Vault. The increased participation directly mitigates volatility and enhances price accuracy, ensuring that RIV Coin functions reliably as a medium of exchange, a numeraire, and a store of value. In essence, the very openness of RIV Coin’s ecosystem enables it to capture the equilibrium advantages outlined in the theoretical results, creating a more transparent, liquid, and resilient market for all participants.
4 Solution
RIV Coin is a cryptocurrency backed by a diversified set of reserves (the Vault) pledged in the exclusive interest of its users. The Vault is actively invested in a diversified CeFi/DeFi solution that does so under transparent, on-chain AND off-chain governance.
The Vault is composed of two interconnected entities:
@@ 4 Solution
    A professionally managed, segregated, compartment of an alternative fund containing real-world assets, whose operational events and state transitions are aggregated using a zero-knowledge Proof of Integrity system, hereinafter called the “Atlas” (see Section 4.1)
@@
    An on-chain reserve holding digital assets directly controlled by only the protocol, hereinafter called the “Aegis” (see Section 4.2.2)
Together, these elements provide a comprehensive, verifiable reserve that anchors the intrinsic value of RIV Coin.
4.1 Atlas Ledger and Proof of Integrity
Atlas is a zk-Rollup specifically designed for investment fund operations, enabling transparent and verifiable fund accounting without revealing sensitive internal data.
Figure 1: Atlas Proof of Integrity System
Definition (zk-Rollups).
Aggregators in zk-Rollups submit a succinct cryptographic proof (called a validity proof) that asserts the correctness of a new published state root relative to the transaction data. This proof guarantees that the transaction execution was exact without requiring the verifier to recompute the operations. Zero-knowledge properties ensure that the necessary inputs remain confidential.[rollups]
Just as zk-Rollups aggregate off-chain transactions into verifiable proofs, Atlas aggregates operational events and state transitions of a traditional asset portfolio—including investments, redemptions, valuations, and accounting adjustments—into cryptographic proofs of integrity.
State Definition.
Each ledger state $S$ consists of:
  • A public homomorphic vector commitment to the full fund state,
    • Instantiated as a Pedersen vector commitment $\mathcal{C}(S)$, satisfying
      $\mathcal{C}(S+S^{\prime})=\mathcal{C}(S)\cdot\mathcal{C}(S^{\prime}),$
      ensuring additive homomorphism.
    • Internally includes:
        A vector of asset quantities,
        A vector of asset prices.
  • A public Net Asset Value (NAV) field stored separately in cleartext after each block.
Asset quantities, prices, and NAV are represented as fixed-point integers over the prime field $\mathbb{F}$, scaled appropriately to prevent rounding errors and modular wraparound. The zk-STARK circuit enforces that, after sequentially applying the batch of transactions, the publicly stored NAV matches the dot product of the committed asset quantities and prices:
$\text{NAV}=\langle\text{assets},\text{prices}\rangle.$
Within the circuit, range constraints ensure all values remain within safe bit-length bounds to avoid wrap-around.
System Design.
Atlas implements a Zero Knowledge Scalable Transparent ARgument of Knowledge (zk-STARK)[starks] over the relation
$R\subseteq\mathcal{X}\times\mathcal{W},$ (41)
where:
  • $\mathcal{X}$ is the set of public inputs,
  • $\mathcal{W}$ is the set of private witnesses,
  • $(x,w)\in R$ if and only if $w$ satisfies the correctness predicate for $x$.
In Atlas:
  • $x=(S_{0},S_{n},h,\mathrm{NAV})$, where:
    • $S_{0}$ is the homomorphic commitment to the initial asset and price vectors,
    • $S_{n}$ is the homomorphic commitment to the final asset and price vectors,
    • $h$ is the Merkle root of the transaction batch,
    • $\mathrm{NAV}$ is the public Net Asset Value after processing the batch.
  • $w=(T,\mathrm{assets},\mathrm{prices})$, where:
    • $T$ is the ordered list of transactions.
    • $\mathrm{assets}$ is the assets vector.
    • $\mathrm{prices}$ is the prices vector.
The verification condition is that there exists a valid witness $w$ such that:
$R(x,w)\iff$ $S_{n}=\mathcal{C}(\mathcal{A}(S_{0},T))$
$\land\ h=\texttt{MerkleRoot}(T)$
$\land\ \text{NAV}=\langle\text{assets},\text{prices}\rangle$ (42)
where:
  • $\mathcal{A}(S_{0},T)$: Sequentially applies the transactions $T$ starting from $S_{0}$
  • $\mathcal{C}(S)$: Computes the cryptographic commitment to a state $S$,
  • $\texttt{MerkleRoot}(T)$: Computes the Merkle root over the batch of ordered transactions.
Only the Merkle root $h$ of transactions $T$ is made public; the full transaction list $T$ remains confidential unless individual transactions are selectively disclosed via Merkle proofs by the committee. All batches are padded to a fixed size (e.g., next power of two) with dummy transactions to avoid metadata leakage.
Transaction Types.
Transactions belong to a predefined set:
$\mathcal{T}=$ ($(\{\texttt{Deposit}\}\times\mathcal{U}\times\mathbb{F})$
[…]
where $\mathcal{U}$ is the set of user addresses and $\mathbb{F}$ is the prime field underpinning the zk-STARK. Each transaction is defined in Algorithm 1:
Algorithm 1 Transaction Definitions
1:  type Deposit
2:      Amount: Fiat amount deposited
3:      User: User ID of the depositor
4:
5:  type Withdrawal
6:      Amount: Fiat amount withdrawn
7:      User: User ID of the withdrawer
8:
9:  type AssetPurchase
10:      Index: Index of the asset purchased in the assets vector
11:      Quantity: Amount of the asset purchased
12:      Price: Effective unit price of the asset at purchase, net of fees.
13:
14:  type AssetSale
15:      Index: Index of the asset purchased in the assets vector
16:      Quantity: Amount of the asset sold
17:      Price: Effective unit price of the asset at sale, net of fees.
18:
19:  type ManagementFee
20:      FeeAmount: Management fee deducted
21:
22:  type PerformanceFee
23:      FeeAmount: Performance fee deducted
24:
25:  type MarkToMarket
26:      NewPrices: Vector of updated asset prices
The MarkToMarket transaction updates the full vector of asset prices at once, where $n$ is the amount of assets, and by extension prices.
Within each block, only the final MarkToMarket transaction is used to update the prices vector for purposes of NAV computation. Intermediate MarkToMarket transactions, if present, are ignored.
Ledger Structure.
Each batch of operational transactions is structured as a block containing:
  • A block number indicating the ledger sequence,
  • The Merkle root $h$ of the ordered transactions included in the block,
  • The old state commitment $S_{0}$ and the new state commitment $S_{n}$,
  • The public NAV value after processing the block.
Each proof $\pi$ associated with a block is co-signed by fund administrators and posted on RIV Chain, where the dedicated x/atlas module verifies it against the relation $R$. State commitments are chained: the new state $S_{n}$ from block $k$ becomes the old state $S_{0}$ for block $k+1$, ensuring sequential integrity.
Security and Transparency.
By combining zk-Rollup architecture with transaction Merkle batching and succinct validity proofs, Atlas achieves:
  • Scalability: Compressing many transactions into a single succinct proof,
  • Privacy: Keeping sensitive financial operations confidential while still verifiable, with transaction batch padding to prevent metadata leakage,
  • Transparency: Providing public, tamper-proof assurance of fund accounting integrity,
  • Selective Auditability: Allowing individual transactions to be selectively revealed via Merkle proofs without exposing unrelated data.
Fund units of the vehicle of investment are entrusted to a Swiss trust governed by a committee of trustees. To ensure coherent oversight, the trustees are the same entities comprising the DAO’s committee (detailed further in Section 5.1), thus aligning on-chain governance with off-chain fiduciary obligations.
At launch, Atlas will focus on investing in a diversified Luxembourg SICAV RAIF (Reserved Alternative Investment Fund) which diversifies into Fixed Income listed securities, thematic equity portfolios (listed stocks), and a diversified number of mutual/hedge/ETFs funds. This provides efficient, capital-light access to a wide range of assets. As Atlas accumulates more capital, it will pursue further diversification by reinvesting into additional asset classes.
4.2 RIV Chain Architecture
RIV Chain is a Layer-1 sovereign blockchain implemented on top of the Cosmos-SDK. It uses a Proof-of-Stake consensus powered by CometBFT: a distributed, Byzantine fault-tolerant, deterministic state machine replication and consensus engine, which is the successor to Tendermint Core [buchman2019latestgossipbftconsensus, kwon2014tendermint]. Figure 2 illustrates the high-level architecture of a Cosmos-SDK-based blockchain, highlighting the separation between consensus and application logic. Communication between these layers is handled by the Application Blockchain Interface (ABCI), a protocol that allows CometBFT to remain agnostic to the specific application state machine.
Figure 2: Cosmos-SDK Blockchain Architecture Overview
4.2.1 Modules
Cosmos SDK modules define the unique functional and stateful properties of each blockchain application. Modules can be viewed as independent state machines operating within the larger application-level state machine. They encapsulate critical elements of blockchain behavior, including storage layouts (state) and state transition functions (message handlers and transactions). Figure 3 illustrates how transactions are routed and handled by their respective modules.
For the purposes of this whitepaper, we will not be exhaustively covering all of the modules used by RIV Chain, such as the core modules and UX-related supplementary modules, but only the ones relevant to RIV Coin. Mainly, that includes the Inter-Blockchain Communication (IBC) modules, and the custom modules implemented by RIV Chain.
Inter-Blockchain Communication (IBC) Modules
RIV Chain leverages the Inter-Blockchain Communication (IBC) protocol to interact with external blockchains and services. IBC allows RIV Chain to perform secure, trust-minimized communication across sovereign chains, enabling a wide range of cross-chain functionalities. The primary IBC modules included are:
  • x/core – Provides the underlying IBC protocol logic, including connection, channel, and packet lifecycle management.
  • x/apps – Implements application-layer protocols built on top of IBC, such as Interchain Accounts (ICA) and Interchain Queries (ICQ).
  • x/light-clients – Manages the verification of counterparty chain states via light clients (e.g., Tendermint-based light clients).
RIV Chain’s use of IBC includes:
  • Interchain Accounts – Allowing RIV Chain to control accounts on remote chains like Osmosis, enabling seamless liquidity management and DeFi interactions without bridging.
  • Osmosis Liquidity Pools – Directly owning and managing liquidity positions in Osmosis pools through Interchain Accounts, enabling native liquidity provision and yield strategies.
  • Band Protocol Oracles – Accessing off-chain price feeds via Band Protocol’s IBC-enabled oracles to obtain reliable real-time asset pricing for portfolio valuation and risk management.
Custom Modules
Last, but not least, the custom modules purpose built for RIV Chain and RIV Coin:
  • x/atlas
  • x/aegis
  • x/emissions
  • x/escrow
x/atlas was explained in Section 4.1, the remaining are explained in the following sections.
Figure 3: Cosmos-SDK Modules Message Processing
4.2.2 Aegis
Aegis refers to a special account on the blockchain that is not governed by private keys. Instead, the assets held within Aegis can only be moved or utilized through operations explicitly defined by the module itself. This design enables the protocol to directly own and manage digital assets without relying on individual custodians. Assets under Aegis may reside natively on the RIV Chain, or, through the Interchain Accounts (ICA) standard, on other IBC-enabled blockchains. At the time of writing, Aegis supports two authorized operations:
  1. 1.
    Transfers to Atlas
  2. 2.
    Protocol-Owned Liquidity (POL) operations (see Section 5.1.1)
For an explanation of how assets are deposited into Aegis, refer to Section 5.2.1.
Aegis lays the foundation for broader autonomous DAO-controlled execution management, with additional capabilities currently under research and planned for future implementation.
4.2.3 Escrow Mechanism
Escrow facilitates secure, auditable transfers of digital assets from Aegis to the fund’s custodied wallet—under the control of the regulated custodian bank.
The process begins with a MsgInitiateEscrowTransfer, where Aegis moves specific assets into a protocol-controlled escrow account. This account grants the DAO committee multisig-based, least-privilege authorization to:
  1. a)
    Execute transfers only to whitelisted addresses (i.e., the fund’s custodied wallet).
  2. b)
    Authorize Private Key Control Procedure (PKCP) transactions required for KYC compliance.
Once assets are released to the custodied wallet, it counts as a deposit or subscription by the DAO committee and the fund units issued are placed in the trustee’s custody. They can be off-ramped to fiat or traded by the invested fund, and are included in ongoing audits managed through Atlas.
Figure 4: Escrow Mechanism Flow
5 Tokenomics
In this section, we describe the Tokenomics of the RIV Coin ecosystem, covering the structure of governance, liquidity management, staking rewards, minting mechanisms, and the emissions system.
5.1 Decentralized Autonomous Organization (DAO)
Governance of the RIV ecosystem is decentralized and coordinated by a DAO structure. Decision-making authority is vested in a governance committee operating through a multisignature wallet (multisig). The committee is responsible for adjusting protocol parameters, enacting policy changes, and stewarding the health of the system, all while adhering to principles of transparency and democratic participation.
5.1.1 Protocol-Owned Liquidity (POL)
Protocol-Owned Liquidity (POL) refers to liquidity positions that are directly controlled by the DAO, rather than by individual liquidity providers. POL ensures that the protocol always maintains deep, reliable liquidity for RIV Coin markets, reducing slippage and improving price stability. Governance, via the committee’s multisig, controls and allocates POL deployments across different pools and venues. The liquidity positions, however, are owned by Aegis, via the Interchain Accounts standard, where the protocol controls accounts on other blockchains and DEXes, and receives the swap fees rewards.
5.1.2 Staking
Staking rewards are derived from a fraction of the net return generated by the Vault’s treasury assets. The reward distribution occurs on a per-epoch basis, where the epoch length and reward fraction are configurable governance parameters. Adjustments to staking parameters are authorized through governance proposals ratified by the committee’s multisig. This ensures that staking remains flexible to changing market conditions while protecting long-term sustainability.
5.2 Minting Flows
The circulating supply of RIV Coin can change through three mechanisms:
  1. 1.
    Inflationary Minting: Minting that occurs independently of any new reserve deposits in the Vault.
  2. 2.
    Non-Inflationary Minting: Minting backed by verifiable reserve deposits into the Vault.
@@ 5.2 Minting Flows
    Burning: Permanent destruction of tokens, reducing circulating supply.
The staking module manages inflationary minting, which distributes minted RIV Coin as rewards to liquidity providers. The committee’s multisig retains discretionary authority to initiate inflationary minting within strict constraints, including a maximum cap of 100% of the total circulating supply enforced at the DAO level.
Non-inflationary minting follows a secured and verifiable flow (detailed in Section 5.2.1), wherein minted RIV Coin corresponds to reserves pledged to the Vault. The result is a quasi-full hedge, where minted tokens are effectively backed always by the value of assets under custody.
The robustness of this reserve-backed design ensures that systemic collapse would require catastrophic external events (e.g. Lehman default, 2022 crisis) or extraordinary fraud (e.g., Madoff, FTX scandals), both of which are mitigated by segregated accounts, multi-layered proofs, and conservative custody policies. Notably, assets within the Vault cannot be rehypothecated or pledged as collateral.
RIV Coin rests upon the three pillars of MPH:
  • Meritocracy: Users who participate in pooling and staking are proportionally rewarded.
  • Pragmatism: Governance decisions and code implementations prioritize user protection and adaptability.
  • Honesty: Transparent, predictable rule sets ensure fair play with no hidden mechanisms.
The combination of advanced proof algorithms, transparent fund segregation, and strict compliance frameworks ensures that price discovery remains free, organic, and legally sound.
5.2.1 The Emissions Manager and Non-Inflationary Minting
The Emissions Manager (EM) is a core component of the DAO responsible for managing controlled expansions of RIV Coin supply into the market through non-inflationary means.
Initially, during the Presale Phase, fiat capital is deposited into the Vault, allowing RIV Coin to be minted at a fixed 1:1 price against reserves. Presale proceeds are used to continue minting new batches, with a portion allocated to bootstrap liquidity pools. After the presale, the protocol transitions into the Open Market Phase.
In the Open Market Phase, a Liquidity Pool (LP) is established where RIV Coin acts as the quote asset. Users can freely trade into and out of RIV Coin and participate as liquidity providers.
The Emissions Manager operates during this phase by monitoring market conditions. Every 3 epochs, it evaluates the premium (market price / backing price) of RIV Coin and if the premium exceeds a configurable minimum premium threshold, the Emissions Manager mints and sells a computed quantity of RIV Coin on the open decentralized exchange in exchange for fresh reserves. This issuance amount is governed by the following equation:
$r_{t}=\sum_{i=0}^{t-1}r_{i}\times b\times\frac{p+1}{P_{\text{min}}+1}$ (44)
Where:
  • $r_{t}$ = Amount of RIV Coin emitted at epoch $t$
  • $b$ = Base emissions rate
  • $p$ = Current premium (market price / backing price)
  • $P_{\text{min}}$ = Minimum premium threshold
The flow of the Emissions Manager system is illustrated in Figure 5.
Figure 5: Minting flow
Due to the premium requirement, this process is slightly deflationary, as the proceeds of the sale are greater than the intrinsic value of the minted coins. This results in an arbitrage that is completely in favor of the users, increasing the intrinsic value of their RIV Coin. Also as a consequence, the market price is closer to the intrinsic value, due to both the price impact of the sale and the deflationary nature of the mint.
A reasonable, fixed part of the total token supply will be allocated to a “Founder’s Wallet” and it will be used also to compensate the team and the most relevant advisors adherent with the best incentives schemes. These tokens will be vested proportionally to the minting activity. The larger the minting the faster will be the premium that Users will release to this Wallet. This wallet will be filled with a maximum cap of 5% of all the token supply. In other words, for every 95 EUR of collateral, the final token minted could be up to 100.
5.2.2 Natural Arbitrage Mechanisms
RIV Coin’s fully backed design positions it as the quote asset in various liquidity pools, effectively anchoring the relative pricing of other tokens. Rather than depending solely on speculative market forces, RIV Coin’s worth stems from the diversified reserves that underlie it. Consequently, whenever a token’s price–expressed in RIV Coin–deviates from that token’s price on other markets, arbitrageurs quickly capitalize on the discrepancy, driving the RIV Coin-denominated price back toward its intrinsic reserve value. This interplay ensures that RIV Coin remains a stable reference, fostering reliable trading pairs and reinforcing its role as a dependable exchange medium.
5.3 Governance
The DAO operates under a mixed governance model where authority is divided between decentralized tokenholders and an executive Committee operating via a multisignature (multi-sig) wallet. Certain critical operational functions fall under the purview of the Committee for efficiency and risk mitigation, while broader strategic development is directed by tokenholder votes.
All RIV Coin possess equal voting rights, with $n$ RIV equating to $n/S$ of the total voting power, where $S$ is the total circulating supply.
5.4 Tokenholder Voting Authority
The following matters are decided through decentralized voting by all RIV Coin holders:
  • Feature Requests: Tokenholders may propose and vote on new feature implementations, expansions, and initiatives beyond the protocol’s existing scope.
  • Sale of Last Resort: Approval of liquidation of the Vault’s assets and token redemption at intrinsic value during systemic crises.
Voting is conducted on a one-token, one-vote basis, with quorum and approval thresholds established in the DAO’s governance documents.
Committee Authority
The Committee serves as the executive arm of the DAO and operates via a multisignature wallet requiring threshold approval for actions. The following matters fall under the Committee’s purview:
  • Minting Authorization: Approval of inflationary and non-inflationary minting operations within protocol-defined frameworks.
  • Protocol-Owned Liquidity (POL) Management: Allocation and reallocation of liquidity deployments.
  • Emergency Response: Execution of protective actions such as trading halts and liquidity withdrawal freezes during crises.
  • Dynamic Parameter Management: Adjustment of staking rates, emission schedules, and premium thresholds.
  • Upgrades and Migrations: Management and authorization of protocol upgrades, infrastructure migrations, and architectural changes.
  • Fund Management Oversight: Appointment and supervision of the external asset managers responsible for the Vault’s investments.
Committee members act as fiduciaries of the protocol and are subject to disclosure obligations and periodic accountability mechanisms.
Reserves Allocation
Portfolio allocation of the Vault’s reserves falls under the purview of the DAO Committee. As detailed in Section 4.1, initial investments focus on centralized, regulated vehicles with a track record of performance. Investment strategies aim for a prudent balance between risk and return, exclusively prioritizing the interests of RIV Coin holders.
Configuration Parameters
The following dynamic parameters are managed by the Committee and may be adjusted periodically based on market conditions and protocol performance:
  • Staking Rate: The fraction of the Vault’s net yield distributed to staking participants per epoch.
  • Base Emission Rate: The fundamental rate at which new RIV Coin is emitted via bond markets, subject to premium conditions.
  • Minimum Premium: The minimum premium (market price relative to backing value) required to trigger new emissions into the market.
These parameters are critical to balancing the protocol’s growth dynamics, ensuring stability, aligning incentives, and maintaining reserve-backed value over time.
Migration Mechanism
The authority to manage migrations and upgrades falls under the purview of the Committee. If necessary, the Committee may authorize migration to new contracts or architectures to enhance protocol functionality, security, or scalability.
Sale of Last Resort
The sale of last resort is a protective mechanism allowing tokenholders to exit the system at intrinsic value during extreme adverse events.
  • Initiation: The DAO Committee may propose activating the sale of last resort.
  • Approval: Requires approval through a tokenholder vote.
  • Execution: Upon approval, the Vault’s assets are liquidated, and proceeds are used to repurchase and burn RIV Coins at the current NAV intrinsic price.
This mechanism ensures that tokenholders retain the right to redeem real value even under distressed market conditions.
5.5 Performance-Based Reserve Management
The fund that Atlas tracks, which constitutes the centralized component of the Vault’s reserves, is subject to standard asset management fees designed to sustain professional oversight, maintain operational excellence, and ensure long-term alignment with tokenholder interests.
The fee structure is composed of two elements:
  • Performance Fees: Fees applied to positive net performance of the fund, designed to align the incentives of the asset managers with the success of the Vault.
    • 0–9% NAV increase (monthly): 10% performance fee
    • 9–20% NAV increase (monthly): 15% performance fee
    • 20%+ NAV increase (monthly): 25% performance fee
    Performance fees are calculated on a monthly basis relative to the Net Asset Value (NAV) growth and are deducted from the gross returns before they are reflected in the Vault’s proofs and reporting.
  • Management Fee: A fixed annualized fee of 2%, charged pro rata and deducted on a monthly basis from the fund’s total assets under management (AUM).
These fees are industry-standard for actively managed portfolios with diversified and sophisticated investment strategies. They compensate professional asset managers for research, portfolio construction, risk management, legal compliance, auditing, and custodial arrangements critical to the sustainable growth of the reserves.
Importantly, all fees are already incorporated into the Net Asset Value (NAV) calculations reflected in Atlas proofs. Users interacting with RIV Coin or referencing the Vault’s performance are always observing figures net of fees, ensuring full transparency and eliminating hidden charges.
The fee structure is designed with the following principles:
  • Alignment of Incentives: Managers are rewarded only when net positive returns are achieved, ensuring their goals align with those of tokenholders.
  • Verifiability: All fee deductions are reflected on-chain through Atlas’ Proof of Integrity and are auditable.
  • Fairness: No performance fee is charged in periods of negative returns, and fees scale progressively with fund performance to avoid excessive extraction.
Through this structure, the fund achieves a balance between incentivizing professional active management and preserving maximum value accrual for RIV Coin holders.